# Amazon Web Service

### v1

@@ 1 Introduction
When asked whether 9.9 is greater than 9.11, Claude Sonnet 4 confidently states that 9.11 is larger, explaining that “since 90 is greater than 11 in the hundredths place, 9.11 is the larger number”—while simultaneously calculating that 9.11 - 9.9 = 0.21 (incorrect). GPT-4o claims that “9.11 is greater than 9.9” in mathematical contexts, invoking software versioning to justify the confusion. Yet when asked to explain how to compare decimal numbers, both models provide flawless algorithmic descriptions: “Write the numbers one above the other, aligning the decimal points… Compare digit by digit from left to right.” Claude Sonnet 4 even works through the exact same comparison correctly as an instructional example, concluding that “9.90 $>$ 9.11.” Both models articulate decimal comparison procedures with textbook precision while failing to execute them reliably.
@@
This disconnect reveals a fundamental limitation in Large Language Models: they exhibit comprehension without competence—a systematic dissociation where models can perfectly explain principles they cannot reliably execute. We term this phenomenon computational split-brain syndrome, drawing analogy to neurological conditions where different brain systems cannot coordinate effectively. Like patients who can verbally describe actions they cannot perform, LLMs develop geometrically separated pathways for “knowing about” procedures versus “executing” them.
@@
This is not a limitation of scale, training data, or prompting techniques. We argue that computational split-brain syndrome stems from structural constraints in how LLMs represent symbols, learn procedures, and execute reasoning steps. This dissociation challenges fundamental assumptions about intelligence derived from human cognition, where explanatory fluency typically correlates with execution competence. LLMs systematically violate this expectation, revealing that artificial and biological intelligence may operate under fundamentally different principles. Unlike symbolic systems that bind tokens to abstract roles and apply rules over those bindings, LLMs operate as pattern completion engines optimized for next-token prediction. This architectural difference creates three interdependent constraints that prevent reliable symbolic execution despite perfect comprehension:
  • Contextual Averaging: Token embeddings encode context-weighted averages that systematically distort symbolic relationships, preventing stable domain binding. The token “9.11” retains historical event associations that interfere with mathematical computation. Such contamination is inevitable in general-purpose models but absent in narrow-domain training.
  • Architectural Impossibility: Feed-forward networks face fundamental constraints that prevent exact symbolic computation through weight configuration alone, forcing hierarchical pattern assembly rather than principled algorithmic implementation.
  • Instruction-Execution Disconnect: Next-token prediction treats algorithmic descriptions and execution traces as equivalent pattern completion tasks, creating geometrically separated representational pathways for procedural knowledge versus computational implementation.
We demonstrate that these constraints manifest consistently across two critical domains: symbolic computation (arithmetic operations) and relational reasoning (logical inference and variable binding). In both domains, models exhibit the same split-brain syndrome—fluent explanation coupled with unreliable execution—revealing this as a fundamental architectural limitation rather than a domain-specific failure.
@@
These architectural limitations manifest consistently across arithmetic and logical domains, creating the systematic dissociation between comprehension and competence we term computational split-brain syndrome.
@@
Our findings suggest that mechanistic interpretability studies identifying “arithmetic circuits” or “adder neurons” may be observing sophisticated forms of pattern matching coordination that are formed path-dependently during learning, rather than genuine computational subroutines. More critically, the geometric dissociation between instruction and execution pathways raises concerns about both model self-explanations and interpretability research: models may articulate reasoning procedures through neural routes distinct from those used for actual computation, potentially misleading both self-monitoring and researcher analysis of neural activations.
These constraints appear unavoidable within the current paradigm: contextual averaging emerges inevitably from next-token prediction over diverse corpora, while predicting next tokens on internet-scale data forces FFNs to work with attention layers into hierarchical pattern assembly rather than principled computation, and instruction-execution disconnect results from treating all text as equivalent prediction targets. This suggests computational split-brain syndrome will persist across model families and scales unless addressed through fundamental architectural innovations rather than incremental improvements.
@@
Section 2 reviews foundational work on transformer computation, symbolic reasoning, and recent interpretability findings that inform our analysis. In Section 3, we investigate LLM failures in symbolic computation, focusing on arithmetic tasks that reveal unstable embeddings and residual fitting behavior. Section 4 extends this analysis to relational reasoning, demonstrating that the same architectural bottlenecks underlie LLM failures in multi-step inference, variable binding, and logical consistency. Section 5 examines compensatory approaches—self-scaffolding, tool delegation, and hybrid architectures—revealing how all three strategies converge on the same reliable metacognitive limitations that current architectures lack. Section 6 analyzes LLMs as hierarchical pattern completion engines, distinguishing between general intelligence (which LLMs achieve through sophisticated pattern matching) and generalizable intelligence (which requires systematic rule discovery and principled reasoning), and examining the performance cliffs that emerge when tasks transition from pattern recognition to genuine algorithmic discovery.
@@
This analysis focuses on pretrained transformer-based LLMs trained on next-token prediction objectives over natural language corpora, without access to external tools or explicit reasoning scaffolds. Our empirical evaluation centers on specific model families (LLaMA2, Claude, GPT-4) and may not generalize to fundamentally different architectures or training paradigms. The geometric separation experiments rely on t-SNE projections which may not capture all relevant representational structure. Our theoretical analysis of FFN computational limits focuses on ReLU networks and may not apply to other activation functions. Additionally, we do not extensively evaluate recent architectural innovations like mixture-of-experts or specialized reasoning modules that might partially address the constraints we identify.
@@ 2.1 Foundations and Theoretical Limits
Recent theoretical frameworks have clarified the fundamental boundaries of transformer expressivity. Strobl et al. (Strobl et al., 2024) provide a comprehensive survey documenting how transformers function as recognizers of formal languages, revealing that while scaffolding strategies can improve performance within bounded expressivity classes, they remain fundamentally constrained by architectural limitations rather than achieving true symbolic reasoning capabilities. The Simons Institute workshop (Simons Institute for the Theory of Computing, 2024) established that transformers act as highly parallel pattern-matching engines with limited capacity for long-range sequential computation, constraining their theoretical expressivity in structured symbolic tasks.
These formal limitations manifest practically as a persistent gap between what LLMs can articulate versus what they can reliably execute. This disconnect becomes especially apparent in arithmetic and logic tasks, where scaffolding strategies like chain-of-thought prompting (Wei et al., 2022) and reasoning-trained models like o1 and DeepSeek-R1 (OpenAI, 2024; DeepSeek-AI et al., 2025) show improved performance yet remain fundamentally compensatory rather than curative—they bypass core architectural bottlenecks rather than resolving them.
@@
Understanding how reasoning capabilities emerge during training reveals critical insights into LLM limitations. Power et al. (Power et al., 2022) demonstrate that late-phase generalization involves fundamental reorganization of internal representations, showing how training dynamics create path-dependent artifacts that appear as reasoning capabilities but reflect statistical regularities rather than systematic algorithms. Tigges et al. (Tigges et al., 2024) extend this analysis to decoder-only LLMs of moderate scale (up to 2.8B parameters), finding that while computational circuits emerge consistently across scale, their specific implementations vary significantly during training, suggesting that apparent “reasoning circuits” may reflect training-specific pattern coordination rather than universal computational principles, though whether this holds for larger models remains an open question. Recent mechanistic analysis of factual learning provides direct evidence for these training-dependent phenomena. Zucchet et al. (2025) demonstrate that transformer knowledge acquisition follows three distinct phases: learning general statistics, a plateau phase where attention circuits develop, followed by individual-specific knowledge acquisition. Critically, they show that adding new factual knowledge rapidly corrupts existing memories stored in feed-forward layers, confirming that apparent knowledge storage reflects fragile statistical coordination rather than stable representation systems.
@@ 2.2 Empirical Failures Across Domains
Systematic evaluation reveals consistent failures in symbolic computation despite surface fluency. Srivastava et al. (Srivastava et al., 2024) quantify sharp performance drops when problem structures deviate from training distributions. Mirzadeh et al. (Mirzadeh et al., 2024) demonstrate through the GSM-Symbolic benchmark that LLMs exhibit noticeable variance when only numerical values change, with performance drops up to 65% when irrelevant clauses are added. This fragility confirms that current LLMs cannot perform genuine logical reasoning and instead replicate reasoning patterns from training data.
Dziri et al. (Dziri et al., 2023) provide comprehensive analysis showing that transformers solve compositional tasks through “linearized subgraph matching”—memorizing computation patterns rather than learning systematic algorithms. Their frequency analysis reveals that models succeed primarily when test computation subgraphs appeared frequently in training data, failing dramatically on out-of-distribution examples. Nikankin et al. (Nikankin et al., 2024) confirm this pattern-dependent behavior, showing that models rely on a “bag of heuristics” for arithmetic rather than implementing systematic algorithms. Wu et al. (Wu et al., 2025) further demonstrate that domain-specific fine-tuning can increase answer accuracy while degrading the logical coherence of chain-of-thought explanations, reinforcing the comprehension–execution split.
Multiple recent studies provide converging evidence for the execution limitation across different domains. Yang et al. (Yang et al., 2024) show that LLMs consistently outperform chain-of-thought when generating Prolog programs for external execution rather than attempting direct calculation. The Illusion of Thinking commentary (Opus and Lawsen, 2025) demonstrates similar improvements when LLMs generate Lua code for Tower of Hanoi problems rather than attempting direct move enumeration. Cheng et al. (Cheng et al., 2024) provide systematic evidence that LLMs excel at pattern recognition and code generation while struggling with deductive reasoning, with their successful framework architecturally separating symbolic code generation from symbolic code execution. This convergent pattern across arithmetic reasoning, logic puzzles, and systematic inference demonstrates that LLMs excel at extracting predicates and generating symbolic representations while requiring external systems for reliable computation—precisely the instruction-execution disconnect we formalize.
@@
The same architectural constraints manifest in relational reasoning. Berglund et al. (Berglund et al., 2023) identify the “reversal curse,” where models fail to infer bidirectional relations from symmetric facts due to asymmetric training exposure. Nezhurina et al. (Nezhurina et al., 2024) show similar failures in multi-step family reasoning problems that should be trivial for systematic logical processing.
@@
Li et al. (Li et al., 2024) provide systematic evaluation across inductive logic programming tasks, finding that LLMs—despite being 100,000 times larger than specialized logical reasoning models—perform dramatically worse on tasks requiring variable binding and systematic rule application. This performance gap persists even when given explicit logical structure through truth-value matrices, definitively ruling out training data scarcity as an explanation and confirming architectural limitations.
@@ 2.3 Mechanistic Explanations
Recent mechanistic interpretability research has developed sophisticated tools for tracing computational pathways within transformer models. Attribution graph studies by Lindsey et al. (Lindsey et al., 2025) demonstrate methods for mapping how models transform inputs to outputs through intermediate representational steps, while other circuit analyses reveal consistency patterns across training and scale (Tigges et al., 2024).
@@
The instruction-execution disconnect operates at the representational level through systematic failures in symbolic binding. McLeish et al. (McLeish et al., 2024) demonstrate that altering input geometry alone can dramatically improve simple numerical reasoning, confirming that embedding representations constitute a significant bottleneck for symbolic manipulation. This supports our analysis that contextual averaging prevents stable domain binding required for symbolic manipulation.
@@
Mechanistic interpretability reveals that apparent “reasoning circuits” often reflect distributed heuristics rather than principled computation. Nikankin et al. (Nikankin et al., 2024) show that arithmetic behaviour emerges from a “bag of heuristics” spread across layers instead of dedicated algorithmic modules. Tigges et al. (Tigges et al., 2024) further demonstrate that, even when a high-level algorithm seems stable, the specific attention heads instantiating it drift throughout training, underscoring strong path dependence. Ye et al. (Ye et al., 2025) train a 124M-parameter GPT-2 exclusively on a procedurally generated grade-school math curriculum and uncover a linearly decodable dependency-graph representation that tracks which intermediate quantities depend on which others. We interpret these coherent circuits as artifacts of an unnaturally homogeneous training path; they dissolve once heterogeneous text is introduced, reinforcing our claim that stable algorithmic modules are fragile in realistic open-domain settings.
@@
Comprehensive analysis of reasoning faithfulness reveals fundamental limitations in how LLMs construct explanations. Plaat et al. (Plaat et al., 2024) survey extensive evidence that LLM reasoning may be “post-hoc” and “constructed after a conclusion has been found,” with larger models showing less faithful reasoning. Chain-of-thought continues to work “even with invalid steps in the reasoning chain,” suggesting pattern matching rather than logical execution. This evidence strongly supports our geometric separation hypothesis—models access instructional pathways distinct from computational pathways when generating explanations.

### v2

@@ 1 Introduction
When asked whether 9.9 is greater than 9.11, Claude Sonnet 4 confidently states that 9.11 is larger, explaining that “since 90 is greater than 11 in the hundredths place, 9.11 is the larger number”—while simultaneously calculating that 9.11 - 9.9 = 0.21 (incorrect). GPT-4o claims that “9.11 is greater than 9.9” in mathematical contexts, invoking software versioning to justify the confusion. Yet when asked to explain how to compare decimal numbers, both models provide flawless algorithmic descriptions: “Write the numbers one above the other, aligning the decimal points… Compare digit by digit from left to right.” Claude Sonnet 4 even works through the exact same comparison correctly as an instructional example, concluding that “9.90 >> 9.11.” Both models articulate decimal comparison procedures with textbook precision while failing to execute them reliably.
@@
This disconnect reveals a fundamental limitation in Large Language Models: they exhibit comprehension without competence—a systematic dissociation where models can perfectly explain principles they cannot reliably execute. We term this phenomenon computational split-brain syndrome [^1] The term also appears in distributed systems to describe partitioned networks that cannot coordinate (Brewer, 2000)—we use the neuroscience analogy as it better captures the functional dissociation between capabilities., drawing analogy to neurological conditions where different brain systems cannot coordinate effectively. Like patients who can verbally describe actions they cannot perform, LLMs develop geometrically separated pathways for “knowing about” procedures versus “executing” them.
Here, comprehension and competence are understood behaviorally—when an LLM flawlessly explains decimal comparison algorithms, observers reasonably judge it as “comprehending” the procedure, just as we assess human understanding through explanation ability. Yet the model cannot reliably execute what it explains. Both capabilities emerge from pattern completion but as geometrically separated, uncoordinated pathways—explanation without execution. This represents an inversion of Dennett’s framework where competence typically precedes comprehension in biological systems (Dennett, 2017).
This split-brain syndrome manifests most visibly as computational hallucination—when models must execute multi-step algorithms they cannot actually perform, they generate plausible-sounding answers through pattern completion rather than computation. Unlike hallucinations that contradict facts or context, computational hallucination emerges directly from architectural impossibility: the model compares 9.11 with 9.9 not through arithmetic but by pattern-matching what such calculations typically look like. Every task requiring symbolic manipulation—from arithmetic to logical inference to data operations—becomes a source of systematic hallucination, with models confidently producing answers that seem reasonable but lack computational grounding.
@@
This is not a limitation of scale, training data, or prompting techniques. We argue that computational split-brain syndrome stems from structural constraints in how LLMs represent symbols, learn procedures, and execute reasoning steps. This dissociation challenges fundamental assumptions about intelligence derived from human cognition, where explanatory fluency typically correlates with execution competence. LLMs systematically violate this expectation, revealing that artificial and biological intelligence may operate under fundamentally different principles. Unlike symbolic systems that bind tokens to abstract roles and apply rules over those bindings, LLMs operate as pattern completion engines optimized for next-token prediction.
@@
These three constraints must be understood together to explain the computational split-brain syndrome:
Claims 1 & 2 (Execution failures): When encountering execution instances (e.g., ‘9.11 > 9.9?’), contextual averaging prevents clean mathematical binding (Claim 1). Even if binding were perfect, transformers face an architectural bottleneck: attention identifies operations and operands but cannot generate new values, forcing FFNs to produce results—yet FFNs cannot implement exact operations, only pattern approximations (Claim 2). Together, these establish that reliable execution is impossible.
Claim 3 (The disconnect): Next-token prediction provides no mechanism for well-learned instructions to guide execution. Both capabilities emerge as separate pattern-matching pathways with no automatic binding between them. The model thus can explain perfectly yet fail at execution—the split-brain syndrome.
Importantly, we demonstrate that these constraints manifest consistently across two critical domains: symbolic computation (arithmetic operations) and relational reasoning (logical inference and variable binding). In both domains, models exhibit the same split-brain syndrome—fluent explanation coupled with unreliable execution—revealing this as limitation of the architecture-training paradigm combination rather than a domain-specific failure.
We now face a fundamental design trade-off. By reducing all language processing to a single continuous optimization task (next-token prediction) over unified token embeddings, transformers achieve their remarkable generality and fluency. However, this same unification prevents the type-based reasoning and domain-specific binding that symbolic computation requires. The contextual averaging that enables rich semantic understanding also contaminates mathematical symbols; the single objective that scales so effectively treats instruction and execution as indistinguishable patterns. Understanding this trade-off clarifies why the computational split-brain syndrome persists across model families and scales: it emerges from the very design choices that make LLMs powerful.
@@
Our findings suggest that mechanistic interpretability studies identifying “arithmetic circuits” or “adder neurons” may be observing sophisticated forms of pattern matching coordination that are formed path-dependently during learning, rather than genuine computational subroutines. More critically, the geometric dissociation between instruction and execution pathways raises concerns about both model self-explanations and interpretability research: models may articulate reasoning procedures through neural routes distinct from those used for actual computation, potentially misleading both self-monitoring and researcher analysis of neural activations. Section 7 extends this analysis to modern interpretability methods, examining whether Sparse Autoencoders and feature-level decomposition escape these limitations or merely surface more stable statistical patterns.
These constraints appear unavoidable within the current paradigm: contextual averaging emerges inevitably from next-token prediction over diverse corpora, while predicting next tokens on internet-scale data forces FFNs to work with attention layers into hierarchical pattern assembly rather than principled computation, and instruction-execution disconnect results from treating all text as equivalent prediction targets. This suggests that computational split-brain syndrome will persist across model families and scales unless addressed through fundamental architectural innovations rather than incremental improvements.
@@
Section 2 reviews foundational work on transformer computation, symbolic reasoning, and recent interpretability findings that inform our analysis. In Section 3, we investigate LLM failures in symbolic computation, focusing on arithmetic tasks that reveal unstable embeddings and residual fitting behavior. Section 4 extends this analysis to relational reasoning, demonstrating that the same architectural bottlenecks underlie LLM failures in multi-step inference, variable binding, and logical consistency. Section 5 examines compensatory approaches—self-scaffolding, tool delegation, and hybrid architectures—revealing how all three strategies converge on the same reliable metacognitive limitations that current architectures lack. Section 6 analyzes LLMs as hierarchical pattern completion engines, distinguishing between general intelligence (which LLMs achieve through sophisticated pattern matching) and generalizable intelligence (which requires systematic rule discovery and principled reasoning), and examining the performance cliffs that emerge when tasks transition from pattern recognition to genuine algorithmic discovery. Section 7 explores how our framework applies to mechanistic interpretability research, examining whether current methods like Sparse Autoencoders identify genuine computational mechanisms or statistical patterns, and proposing tests to distinguish between these possibilities. Section 8 offers a brief reflection on the research journey that led to these insights.
@@
This analysis focuses on pretrained transformer-based LLMs trained on next-token prediction objectives over natural language corpora, without access to external tools or explicit reasoning scaffolds. We analyze these “pure” systems to understand why compensatory strategies have become essential—revealing not just what fails, but why tool use, scaffolding, and hybrid architectures represent necessary workarounds rather than optional enhancements for fundamental architectural constraints. Our empirical evaluation centers on specific model families (LLaMA2, Claude, GPT-4) and may not generalize to fundamentally different architectures or training paradigms. The geometric separation experiments rely on t-SNE projections which may not capture all relevant representational structure. Our theoretical analysis of FFN computational limits focuses on ReLU networks and may not apply to other activation functions. While recent architectural variations like mixture-of-experts or reasoning-augmented models may improve performance, they operate within the same fundamental transformer framework and thus likely face similar constraints—though empirical validation of this hypothesis remains future work.
@@ 2.1 Foundations and Theoretical Limits
Recent theoretical frameworks have clarified the fundamental boundaries of transformer expressivity. Peng et al. (2024) use communication complexity to prove that transformer layers cannot compose functions (e.g., identifying grandparents in genealogies) when domains are sufficiently large. Strobl et al. (2024) provide a comprehensive survey documenting how transformers function as recognizers of formal languages, revealing that while scaffolding strategies can improve performance within bounded expressivity classes, they remain fundamentally constrained by architectural limitations rather than achieving true symbolic reasoning capabilities. The Simons Institute workshop (Simons Institute for the Theory of Computing, 2024) established that transformers act as highly parallel pattern matching engines with limited capacity for long-range sequential computation, constraining their theoretical expressivity in structured symbolic tasks.
However, Li et al. (2024b) prove that scaffolding strategies, techniques in which models generate intermediate steps to guide reasoning, can theoretically expand computational power: While constant-depth transformers can only solve problems in $\mathsf{TC}^{0}$, adding $T$ intermediate steps enables them to solve any problem solvable by Boolean circuits of size $T$. Chain-of-thought prompting (Wei et al., 2022) exemplifies this approach: instead of directly computing ‘23 × 17 = ?’, models generate ‘23 × 10 = 230, 23 × 7 = 161, 230 + 161 = 391’, converting parallel constraints into serial computation. These analyses can be seen as the theoretical foundation for the use of self-scaffolding as a principled workaround (see Section 5). Recent large reasoning models such as o1 and DeepSeek-R1 (OpenAI, 2024; DeepSeek-AI et al., 2025) show improved performance by essentially scaffolding their thought processes at inference time. Yet, they remain fundamentally compensatory rather than curative—they bypass core architectural bottlenecks rather than resolving them.
@@
Understanding how reasoning capabilities emerge during training reveals critical insights into LLM limitations. Power et al. (2022) demonstrate that late-phase generalization involves a fundamental reorganization of internal representations, showing how training dynamics create path-dependent artifacts that appear as reasoning capabilities but reflect statistical regularities rather than systematic algorithms. Tigges et al. (2024) extend this analysis to moderate scale LLMs with decoder only (up to 2.8B parameters), finding that while computational circuits emerge consistently across scale, their specific implementations vary significantly during training. We interpret these findings as suggesting that apparent “reasoning circuits” may reflect training-specific pattern coordination rather than universal computational principles, although this inference requires further validation through controlled experiments with different training permutations.
Recent mechanistic analysis of factual learning provides direct evidence for these training-dependent phenomena. Zucchet et al. (2025) demonstrate that the acquisition of knowledge in transformers follows three distinct phases: general statistics learning, a plateau phase where attention circuits develop, followed by acquisition of individual-specific knowledge. Critically, they show that adding new factual knowledge rapidly corrupts existing memories stored in feedforward layers, confirming that apparent knowledge storage reflects fragile statistical coordination rather than stable representation systems.
@@ 2.2 Empirical Failures Across Domains
Systematic evaluation reveals consistent failures in symbolic computation despite surface fluency. Srivastava et al. (2024) quantify sharp performance drops when problem structures deviate from training distributions. Mirzadeh et al. (2024) demonstrate through the GSM-Symbolic benchmark that LLMs exhibit noticeable variance when only numerical values change, with performance drops up to 65% when irrelevant clauses are added. This fragility confirms that current LLMs cannot perform genuine logical reasoning and instead replicate reasoning patterns from training data.
Dziri et al. (2023) provide comprehensive analysis showing that transformers solve compositional tasks through “linearized subgraph matching”—memorizing computation patterns rather than learning systematic algorithms. Their frequency analysis reveals that models succeed primarily when test computation subgraphs appeared frequently in training data, failing dramatically on out-of-distribution examples. Nikankin et al. (2024) confirm this pattern-dependent behavior, showing that models rely on a “bag of heuristics” for arithmetic rather than implementing systematic algorithms.Wu et al. (2025) further demonstrate that domain-specific fine-tuning can increase answer accuracy while degrading the logical coherence of chain-of-thought explanations, reinforcing the comprehension–execution split.
Multiple recent studies provide converging evidence for the execution limitation across different domains. Yang et al. (2024) show that LLMs consistently outperform chain-of-thought when generating Prolog programs for external execution rather than attempting direct calculation. The Illusion of Thinking commentary (Opus and Lawsen, 2025) demonstrates similar improvements when LLMs generate Lua code for Tower of Hanoi problems rather than attempting direct move enumeration. Wolff and Hulsebos (2025) reveal significant deficits when LLMs attempt tabular reasoning tasks requiring multi-step algorithms such as computing averages or finding maximum values—operations that would be trivial for database systems. Cheng et al. (2024) provide systematic evidence that LLMs excel at pattern recognition and code generation while struggling with deductive reasoning, with their successful framework architecturally separating symbolic code generation from symbolic code execution. This convergent pattern across arithmetic reasoning, logic puzzles, tabular data operations, and systematic inference demonstrates that LLMs excel at extracting predicates and generating symbolic representations while requiring external systems for reliable computation, precisely the instruction-execution disconnect we formalize.
@@
The same architectural constraints manifest in relational reasoning. Berglund et al. (2023) identify the “reversal curse,” where models fail to infer bidirectional relations from symmetric facts due to asymmetric training exposure. Nezhurina et al. (2024) show similar failures in multi-step family reasoning problems that should be trivial for systematic logical processing.
@@
Li et al. (2024a) provide systematic evaluation across inductive logic programming tasks, finding that LLMs—despite being 100,000 times larger than specialized logical reasoning models—perform dramatically worse on tasks requiring variable binding and systematic rule application. This performance gap persists even when given explicit logical structure through truth-value matrices, definitively ruling out training data scarcity as an explanation and confirming architectural limitations.
@@ 2.3 Mechanistic Explanations
Recent mechanistic interpretability research has developed sophisticated tools for tracing computational pathways within transformer models. Attribution graph studies by Lindsey et al. Lindsey et al. (2025b) demonstrate methods for mapping how models transform inputs to outputs through intermediate representational steps, while other circuit analyses reveal consistency patterns across training and scale (Tigges et al., 2024).
@@
The instruction-execution disconnect operates at the representational level through systematic failures in symbolic binding. McLeish et al. (2024) demonstrate that altering input geometry alone can dramatically improve simple numerical reasoning, confirming that embedding representations constitute a significant bottleneck for symbolic manipulation. This supports our analysis that contextual averaging prevents stable domain binding required for symbolic manipulation.
@@
Mechanistic interpretability reveals that apparent “reasoning circuits” often reflect distributed heuristics rather than principled computation. Nikankin et al. (2024) show that arithmetic behaviour emerges from a “bag of heuristics” spread across layers instead of dedicated algorithmic modules. Tigges et al. (2024) further demonstrate that, even when a high-level algorithm seems stable, the specific attention heads instantiating it drift throughout training, underscoring strong path dependence. Ye et al. (2025) train a 124M-parameter GPT-2 exclusively on a procedurally generated grade-school math curriculum and uncover a linearly decodable dependency-graph representation that tracks which intermediate quantities depend on which others. We interpret these coherent circuits as artifacts of an unnaturally homogeneous training path; they dissolve once heterogeneous text is introduced, reinforcing our claim that stable algorithmic modules are fragile in realistic open-domain settings.
@@
Comprehensive analysis of reasoning faithfulness reveals fundamental limitations in how LLMs construct explanations. Plaat et al. (2024) survey extensive evidence that LLM reasoning may be “post-hoc” and “constructed after a conclusion has been found,” with larger models showing less faithful reasoning. Chain-of-thought continues to work “even with invalid steps in the reasoning chain,” suggesting pattern matching rather than logical execution. This evidence strongly supports our geometric separation hypothesis—models access instructional pathways distinct from computational pathways when generating explanations.
Sparse Autoencoders and Feature-Level Analysis.
Recent advances in mechanistic interpretability have shifted from neuron-level analysis to feature-level decomposition through Sparse Autoencoders (SAEs) (Bricken et al., 2023; Templeton et al., 2024). These methods extract interpretable features that appear more stable across models than individual neurons (Wang et al., 2025), with techniques like Linear Computation Graphs tracing complete computational pathways through these features (He et al., 2024). While SAEs promise to overcome the polysemanticity and path-dependency limitations of neuron-level analysis, we examine in Section 7 whether these features represent genuine computational mechanisms or merely more stable statistical patterns. Our analysis suggests that even sophisticated feature decomposition cannot escape the fundamental architectural constraints that create the computational split-brain syndrome.